OpenStudy (anonymous):

3. A train running between two towns arrives at its destination 10 minutes late when it goes 40 miles per hour and 16 minutes late when it goes 30 miles per hour. The distance between the two towns is A. 720 miles B. 12 miles C. 8-6/7 miles D. 12-7/7 miles E. None of these

OpenStudy (anonymous):

Please explain as well .

OpenStudy (zpupster):

distance = rate(time) or d=rt so at 40 mph d = 40(t+10/60) this is in hrs because of mph and we have time plus 10 minutes for 30 mph d= 30(t+16/60) the distance is equal so we set them equal to each other like this: 40t + 400t/60 = 30t + 480/60 we multiplied thru you should be able to get it from here.

OpenStudy (zpupster):

if you cant ill help

OpenStudy (anonymous):

To be honest. Its confusing brother ^^'

OpenStudy (zpupster):

ok you get the d=rt